A 2.0m solution has a boiling point of 103.1oC. The Kb for water is 0.52oC/m, which of the following could be the solute?

a. NaCl
b. KOH
c. CAF2
d. Al2O3
1
Expert's answer
2020-04-27T01:45:10-0400
"\delta T_b= i\times K_b\times m\\\\or \\103.1-100=i\times0.52\times 2\\or\\i=2.98 \ or\ 3"

Dissociation constant = 3

It means 3 ions is released in solution .

So 3 ions is released in the solution

Hence it may be "CaF_2"